Articles on Serre’s Conjectures

1. On modular representations of Gal(\overline{Q}/Q) arising from modular forms. Invent. Math. 100, 431–476 (1990)

2. Bimodules and Abelian Surfaces. Advanced Studies in Pure Mathematics 17, 359–407 (1989)

3. On the Component Groups and the Shimura Subgroup of J_{o}(N). Sém. Th. des Nombres de l’Université de Bordeaux, 1987–88, exposé 6

4. (with B. Mazur) Two-dimensional representations in the arithmetic of modular curves. Astérisque 196–97, 215–255 (1991)

5. Raising the levels of modular representations. Séminaire de Théorie des Nombres, Paris, 1987–88. Progress in Mathematics 81, 259–271 (1990)

6. From the Taniyama-Shimura Conjecture to Fermat’s Last Theorem. Annales de la Faculté des Sciences de l’Université de Toulouse 11, 116–139 (1990)

7. Multiplicities of Galois representations in Jacobians of Shimura curves. Israel Mathematical Conference Proceedings 3, 221–236 (1990)

8. The old subvariety of J_{o}(pq). In: Arithmetic Algebraic Geometry, G. van der Geer, F. Oort, J. Steenbrink, eds. Progress in Mathematics 89, 293– 307 (1990)

9. (with N. Boston and H.W. Lenstra, Jr.) Quotients of group rings arising from two-dimensional representations. C. R. Acad. Sci. Paris Sér. I Math. 312, 323–328 (1991)

10. Multiplicities of p-finite mod p Galois representations in J_{o}(Np). Boletim da Sociedade Brasileira de Matemática, Nova Série, 21, 177–188 (1991)

11. Lowering the levels of modular representations without multiplicity one. International Mathematics Research Notices, 15–19 (1991)

12. Report on mod ℓ representations of Gal(\overline{Q}/Q). Proceedings of Symposia in Pure Mathematics, to appear

K. Ribet